### v1

# Critical spin fluctuations across the superconducting dome in La2-xSrxCuO4

Abstract
Overdoped cuprate superconductors are strange metals above their superconducting transition temperature. In such materials, the electrical resistivity has a strong linear dependence on temperature ($T$) and electrical current is not carried by electron quasiparticles as in conventional metals. Here we demonstrate that the strange metal behavior co-exists with strongly temperature-dependent critical spin fluctuations showing dynamical scaling across the cuprate phase diagram. Our neutron scattering observations and the strange metal behavior are consistent with a spin density wave quantum phase transition in a metal with spatial disorder in the tuning parameter. Numerical computations on a Yukawa-Sachdev-Ye-Kitaev model yield an extended ‘Griffiths phase’ with scaling properties in agreement with observations, establishing that low-energy spin excitations and spatial disorder are central to the strange metal behavior.
Understanding high-temperature superconductivity in layered cuprates has posed significant challenges to the theory of condensed matter. In particular, the unusual metallic state of cuprates above their critical temperature, $T_{c}$, remains poorly understood. This state is crucial, as superconductivity emerges from it. The cuprate phase diagram is characterized by the presence of various competing orders such as antiferromagnetism (AF), charge density wave (CDW) order, and the pseudogap state in the underdoped region of the phase diagram Keimer et al. (2015). The overdoped region is largely free from these competing phases. Nevertheless, the normal state ($T>T_{c}$) shows “strange metal” (SM) behavior Phillips et al. (2023) where the resistivity, $\rho$, is proportional to temperature, $T$, with a large proportionally constant. The resistivity of a metal is usually interpreted using the Drude model in which $\rho\propto\tau_{\text{tr}}^{-1}$, where $\tau_{\text{tr}}^{-1}$ is the relaxation rate of the electrons transporting current. For SMs, we find that $\tau_{\text{tr}}^{-1}\approx k_{B}T/\hbar$, the Planckian dissipative rate Zaanen (2004). This is in stark contrast to the $\rho\propto T^{2}$ and $\tau_{\text{tr}}^{-1}\propto T^{2}$ expected to describe metals in Fermi liquid theory at low temperature. The strange behavior Hartnoll and Mackenzie (2022) in resistivity is accompanied by an anomalous specific heat contribution $\propto T\ln(1/T)$ in contrast to the liquid Fermi form $\propto T$.
The existence of the Planckian transport relaxation rate $\tau_{\text{tr}}^{-1}(T)$ over a wide temperature range in cuprates is challenging for theory to explain. The picture in simple metals is based on the scattering of the fermion quasiparticles near the Fermi surface off bosons (e.g. phonons). The total momentum of the system is then degraded via impurities or umklapp scattering. There are two issues with is picture with regards to cuprates Patel et al. (2023). First, linear or nearly-linear resistivity exists over a wide range of doping so the relevant excitations would also be present over this range. Second, the degradation of momentum mentioned above cannot be accounted for by umklapp scattering. Thus other mechanisms are required.
Here we show that nearly-critical low-energy collective spin fluctuations are integral to the strange metal behavior in the cuprates and exist across the superconducting dome. We investigate the overdoped region a cuprate superconductor where the SM behavior is clearest Cooper et al. (2009). The material we have chosen is La2-xSrxCuO4 (LSCO). This system can be doped over a wide range through the overdoped region of the phase diagram (Fig. 1A) and has a relatively low $T_{c}$ so that the SM normal state is present over a wide range of temperature.
Figure 1: Spin fluctuations in La2-xSrxCuO4 (x=0.22). (A) Schematic phase diagram of LSCO showing the critical spin fluctuations observed here. (B) Schematic of spin excitations in LSCO($x=0.22$) based on (Refs. Lipscombe et al. (2007); Robarts et al. (2019)) as represented by imaginary part of the susceptibility $\chi^{\prime\prime}(\mathbf{Q},\omega)$. Reciprocal space is labeled as $\mathbf{Q}=H\mathbf{a}^{\star}+K\mathbf{b}^{\star}+L\mathbf{c}^{\star}\equiv(H,K)$. Gray region is broad high energy spin excitations Lipscombe et al. (2007); Robarts et al. (2019). Pink regions are the low-energy excitations studied here. (C-F) Slices of $\chi^{\prime\prime}(\mathbf{Q},\omega)$ for $\hbar\omega=4$ meV, $L\in[-1,1]$ and various temperatures (seeSupplementary Sup for details). White dashed lines indicate the range over which slices are integrated to produce cuts in Fig. 2.
Figure 2: The temperature dependence to the low-energy spin fluctuations in La2-xSrxCuO4 (x=0.22). (A) The scattering intensity versus wavevector $\mathbf{Q}$ for different temperatures (bottom to top: 26, 40, 80, 155, 300 K) for energy transfers $\hbar\omega=4,8$ meV. The cuts are through the incommensurate $\mathbf{Q}_{\delta}$ positions as shown by the dashed lines in Fig. 1D. (B) $E-\mathbf{Q}$ map of the magnetic response function $\chi^{\prime\prime}(\mathbf{Q},\omega)$ for $T=26,80$ K showing its evolution with temperature. The trajectory of $\mathbf{Q}$ is the same as that in (A).
Figure 3: Scaling behavior of the dynamic magnetic susceptibility and correlation length in La2-xSrxCuO4 (A) Temperature dependence of $\chi^{\prime\prime}(\mathbf{Q},\omega)$ at ordering wavevector $\mathbf{Q}_{\delta}$. Lines are fits to Eqn. 3. Inset shows the $T$-dependence of the spin relaxation rate $\Gamma_{\delta}$ and real part of the susceptibility $\chi^{\prime}(\mathbf{Q}_{\delta})$ obtained when data are modeled with Eqns. 3-4. (B) $\omega/T$ scaling plot of the $x=0.22$ data in (A) using Eqn. 2 and the procedure described in Sup . The inset shows how the quality of the collapse Sup varies with exponent $\alpha$. (C) Scaling plot for underdoped LSCO $x=0.14$ data from Ref. Aeppli et al. (1997). This data shows a similar collapse to the $x=0.22$ data in (B). (D) The width in $\mathbf{Q}$ of $\chi^{\prime\prime}(\mathbf{Q,\omega})$ is denoted by the inverse dynamic correlation length $\kappa(\omega)$. Here we plot $\kappa(\omega)$ versus temperature for $\hbar\omega=1.25,2.25,4,8$ meV. (E) Data in (D) can be collapsed onto a single trend line with suitable choice of $z$ and $\omega/T$ scaling. (F) Data in (D) can be scaled onto a single line using (7) allowing the dynamic critical exponent $z$ to be determined. Dashed line is Eqn. S15 Sup .
[…]
We used inelastic neutron scattering to map about the $\mathbf{Q}$-$\omega$ dependence of the low-energy spin fluctuation in single crystals of LSCO(x$=0.22$) in the normal state for temperatures $T_{c}\leq T\leq 300$ K. Measurement were carried out using the LET spectrometer of the ISIS Pulsed Neutron and Muon Source Sup . Fig. 1C-F show constant energy maps of the dynamic susceptibility $\chi^{\prime\prime}(\mathbf{Q},\omega)$ at $\hbar\omega=4$ meV. For $T=26$ K, the four-peaks around $\mathbf{Q}=(1/2,1/2)$ can be clearly seen in Fig. 1C. Cuts through the peaks along the $\mathbf{Q}=(1/2-\delta/2+\xi,1/2-\delta/2+\xi)$ line (see dashed lines in Fig. 1D) are shown in Fig. 2A. As the temperature increases, the peaks can be seen to weaken and broaden in $\mathbf{Q}$. The measured intensity $S(\mathbf{Q},\omega)$ can be converted to the the magnetic response function $\chi^{\prime\prime}(\mathbf{Q},\omega)$ using the fluctuation dissipation theorem $S(\mathbf{Q},\omega)=(1/\pi)\chi^{\prime\prime}(\mathbf{Q},\omega)[1-\exp(-\hbar\omega/k_{B}T)]^{-1}$ Sup , the results shown in Fig. 2B and Fig. 3A. Our data (see Fig. 3A) show that the dynamic spin susceptibility near $\mathbf{Q}_{\delta}$ is strongly temperature dependent for this over-doped cuprate. The behavior is reminiscent of the approach to a quantum critical point as $T\rightarrow 0$.
At a quantum critical point, the dynamical susceptibility is expected to show the scaling behavior Sup ; Sachdev and Ye (1992)
$\chi(q,\omega)\propto T^{-\alpha}\Phi\left(\xi q,\frac{\hbar\omega}{k_{B}T}\right),$ (1)
where $\xi\propto T^{-1/z}$ is a correlation length, $z$ is the dynamic critical exponent, $q=\left|\mathbf{Q}-\mathbf{Q}_{\delta}\right|$ is wavevector relative to ordering wavevector, and $\Phi$ is a universal complex function of both arguments. For $q=0$, the imaginary part of the dynamic susceptibility (measured here) scales Sup as
$\chi^{\prime\prime}(q=0,\omega)=T^{-\alpha}\phi_{1}\left(\frac{\hbar\omega}{k_{B}T}\right),$ (2)
where $\alpha=\gamma/\nu z$, $\phi_{1}(x)=\Phi(0,x)$ is a scaling function and $\gamma$ and $\nu$ are the susceptibility and correlation function temperature exponents.
In Fig. 3B we show that our data can be collapsed onto a single trend with a suitable choice of $\alpha$, where $\phi_{1}(x)$ can be approximated by a simple Lorentzian $\phi_{1}(x)\propto ax/(a^{2}+x^{2})$ with $a=2.9\pm 0.3$ and $\alpha=0.32\pm 0.10$. This small value of $\alpha$ excludes the familiar paramagnon theory which has $\alpha=1$ Millis et al. (1990) (which also does not have an extended regime of criticality). Our data is consistent with the susceptibility having a low-frequency component that varies as
$\chi^{\prime\prime}(\mathbf{Q}_{\delta},\omega)=\frac{\chi^{\prime}(\mathbf{Q}_{\delta})\Gamma_{\delta}\,\omega}{\Gamma_{\delta}^{2}+\omega^{2}},$ (3)
where the real part of the susceptibility $\chi^{\prime}(\mathbf{Q}_{\delta})\propto T^{-\alpha}$ and spin relaxation rate varies as
$\hbar\Gamma_{\delta}=ak_{B}T,$ (4)
as shown inset of Fig. 3A. Thus our data is consistent with $\Gamma_{\delta}$ tracking the Planckian relaxation rate $\tau^{-1}_{\text{tr}}\propto T$ as observed in transport measurements suggesting that they are related.
We can gain more information about the criticality of the spin fluctuations from the $\mathbf{Q}$-dependence of $\chi^{\prime\prime}(\mathbf{Q},\omega)$. From Eqn. 3 we see that $\chi^{\prime\prime}(q,\omega)\propto\omega$ as $\omega\rightarrow 0$ so there will be peaks in $\chi^{\prime\prime}(\mathbf{Q},\omega)/\omega$ at the lowest energies. It can be shown from the scaling of the susceptibility Sup that
$\left.\frac{\chi^{\prime\prime}(q,\omega)}{\omega}\right|_{\omega\rightarrow 0}\propto T^{-\left(\alpha+1\right)}\phi_{2}(\xi q).$ (5)
The function $\phi_{2}(\xi q)$ is peaked at the ordering wavevector $q=0$ with width $\xi^{-1}$. We can determine $z$ from the temperature dependence of $\xi^{-1}$.
In order to fit our data, we use a form previously used Aeppli et al. (1997); Zhu et al. (2023) to describe the low-energy response of the cuprates:
$\chi^{\prime\prime}(\mathbf{Q},\omega)$ $=\chi^{\prime\prime}(\mathbf{Q}_{\delta},\omega)\frac{\kappa^{4}(\omega)}{[\kappa^{2}(\omega)+R(\mathbf{Q})]^{2}},$ (6)
where $R(\mathbf{Q})$ is a function with zeros at the $\mathbf{Q}_{\delta}$ positions and the property $R(\mathbf{Q}_{\delta})=|\mathbf{q}|^{2}$ for $\mathbf{q}\rightarrow 0$, the second factor represents $\phi_{2}(\xi q)$ in the limit $\omega\rightarrow 0$ and we can make the association $\kappa(\omega=0)\equiv\xi^{-1}$ Sup . Example fits to Eqn. 6 are shown in Fig. 2. The resulting fitted values of $\kappa(\omega)$ ($\mathbf{Q}$-width of response) are shown in Fig. 3D,E,F. In Fig. 3D we see that $\kappa(\omega)$ increases with $\omega$ and $T$. We posit that the effects of $\omega$ and $T$ add in quadrature and this suggests the form
$\kappa^{2}(\omega)\propto(\hbar\omega)^{2/z}+r(k_{B}T)^{2/z},$ (7)
where $\kappa(\omega=0)=\xi^{-1}\propto T^{1/z}$ and $r$ is the ratio of the temperature and energy contributions which we set equal to one. We find that Eqn. 7 can be used to collapse the experimentally determined $\kappa(\omega,T)$ onto a single solid line as shown in Fig. 3E,F with $z=1.83\pm 0.35$. In the Supplementary Sup we show that our data are also consistent with the system being tuned slightly away from quantum criticality (dashed line in Fig. 3F).
Here we only study the normal state, however, we can extrapolate the data to lower temperatures. Fig. 3 suggests that the spin fluctuations would continue to evolve as $T\to 0$ in the absence of superconductivity. This is supported by measurements made where superconductivity is suppressed by magnetic field. Neutron scattering measurements Zhu et al. (2023) show that $\Gamma_{\delta}$ continues to decrease below $T_{c}$ when a 9 Tesla field is applied. High field ($B\gtrsim B_{c2}$) heat capacity measurements show a $\sim T\log(1/T)$ contribution Girod et al. (2021) and, resistivity measurements in a high field Cooper et al. (2009) show a $\rho\propto T$ behavior as $T\to 0$.
[…]
We now discuss our results in the context of the $p$-$T$ phase diagram (Fig. 1A). A region of glassy spin-freezing is observed Suzuki et al. (1998); Chang et al. (2008); Frachet et al. (2020); Campbell et al. (2024) near $p\approx 1/8$ (also seen in other cuprates Wu et al. (2013)). This region appears to be connected to charge ordering is present in other cuprates Wu et al. (2013) and can be expanded by the application of a large magnetic field Chang et al. (2008); Frachet et al. (2020); Campbell et al. (2024). Inside this region $T$-linear resistivity is not observed Campbell et al. (2024). For larger dopings i.e. $p\gtrsim 0.14$ there is a region where $\omega/T$ scaling in $\chi^{\prime\prime}(\mathbf{Q}_{\delta},\omega)$ and approximately $T$-linear behavior in the resistivity Cooper et al. (2009) co-exist in the absence of spin-freezing.
Understanding the strange metal behavior seen in LSCO and other cuprates is one of the major challenges in condensed matter physics. It occurs across a diverse set of materials that include transition metal oxides, heavy fermions Aronson et al. (1995); Gannon et al. (2018), iron-based superconductors and twisted bilayer graphene Hartnoll and Mackenzie (2022); Phillips et al. (2023). Planckian dissipation provides a universal phenomenology representing a quantum limit on the current-carrying degrees of freedom of a metal Zaanen (2004). However, it does not indicate a microscopic mechanism. The behavior in the resistivity has often been associated with an underlying magnetic quantum critical point (QCP) at $T=0$. In a typical scenario the system is tuned to the QCP and strange metal behaviour is observed over a relatively small range of parameter space. This is observed in some systems such as Sr3Ru2O7 Bruin et al. (2013); Lester et al. (2021). However, there are many examples including not only cuprates but twisted-layer graphene and twisted transitional-metal dichalcogenides where it persists over of an extended region. Extended critical behaviour has also been observed in MnSi with $\rho\propto T^{3/2}$.
Figure 4: Numerical theoretical results for low energy spin fluctuations. (A-D) Compare with the experimental plots in Fig. 3A,B,D,E ($\beta=1/T$). Obtained from the results of Ref. Patel et al. (2024a) (with a mean-field treatment of interactions, but exact treatment of disorder), at the critical value of $\lambda=\lambda_{c}=-0.4586$, after exact analytic continuation to real frequencies Sup ; results for $\lambda>\lambda_{c}$, and parameter values are in the Supplementary Sup , where we also estimate $\approx 33$ meV as the unit of energy/frequency/temperature in the numerics. (A-B) Dynamic spin susceptibility and its scaling plot, with best-fit exponent $\alpha=0.84$. (C-D) Inverse correlation length $\kappa$ and its scaling plot, with best-fit exponent $z=2.128$. (E) Computations of imaginary frequency ($\Omega_{m}$) susceptibility by (in principle, exact) quantum Monte Carlo results for the YSYK model in Ref. Patel et al. (2024b), at the end of the strange metal phase that is present for $\lambda_{s}=5.0\leq\lambda\leq\lambda_{G}=5.5$. The dashed line in the inset is a fit to $\sim T^{\alpha}$. Both plots yield $\alpha=0.20$. (F) Resistivity from spin fluctuations. Computed from the spin susceptibility, as described in Ref. Patel et al. (2024a). Note the distinct slopes of linear-in-$T$ resistivity at low and high $T$.
In recent years new insights have been gained by considering the Hertz-Millis theory Hertz (1976); Millis (1993) for the onset of magnetic order in metals, but in the presence of spatial disorder in the tuning parameter. Such physics can be described by Yukawa-Sachdev-Ye-Kitaev models Esterlis et al. (2021); Patel et al. (2023, 2024a); Li et al. (2024); Patel et al. (2024b), as discussed in the Supplementary Sup . In particular, it has been shown that the inclusion of disorder can explain the existence of the SM-behavior down to the lowest temperatures and over an extended range of doping. Applying this general picture to the cuprates it is postulated that disorder delays the onset of long range spin order giving rise to a “quantum Griffiths phase” Hoyos et al. (2007); Patel et al. (2024a, b) (similar to that in the two-dimensional Ising model in a transverse field Motrunich et al. (2000)) as a transition tuning parameter $\lambda$ is reduced. SM linear-$T$ resistivity is observed over an extended range of $\lambda$ that comprises part of the Griffiths phase instead of QCP. The tuning parameter $\lambda$ is controlled by the doping $x$. The inclusion of disorder is logical in the La2-xSrxCuO4 system because of the perturbation caused by Sr doping in the plane neighboring the CuO2 plane.
Returning to the strange metal behavior in the resistivity observed in LSCO and other cuprates. An extended Drude analysis of optical conductivity $\sigma(\omega)$ of cuprates shows that the optical scattering rate $\tau_{\text{op}}^{-1}(\omega)$ is strongly frequency dependent Carbotte et al. (2011); Michon et al. (2023) indicting the current carrying electrons are scattered inelastically. This is incompatible with elastic scattering from impurities or defects. Phonon scattering cannot produce the $T$-linear resistivity at low temperature below the Debye temperature scales. This has led some to conclude that disorder in strange metals must scatter electrons inelastically Patel et al. (2023); Li et al. (2024). The optical scattering rate $\tau_{\text{op}}^{-1}(\omega)$ provides a further insight into the nature of strange metals as it found to show $\omega/T$ scaling over a wide range of $\omega/T$ Michon et al. (2023).
In Fig. 4 we show numerical calculations of the critical spin susceptibility $\chi^{\prime\prime}(\mathbf{Q}_{\delta},\omega)$ from a Yukawa-Sachdev-Ye-Kitaev model Sup . An approximate method Patel et al. (2024a) allows exact analytic continuation to real frequency, while imaginary time quantum Monte Carlo studies Patel et al. (2024b) yield precise values of exponents. The form (6) is an excellent fit to the theory Sup , and this allows determination of $\kappa(\omega,T)$, which is also shown. The results have been scaled in the same way as Fig. 3B,E and we find excellent real frequency scaling at the quantum critical value of the tuning parameter $\lambda=\lambda_{c}=-0.4586$ with $\alpha=0.84$ and $z=2.13$. Corresponding scaling plots within the quantum Griffiths phase, $\lambda>\lambda_{c}$ appear in the Supplementary Sup : the $\omega/T$ scaling is still present, but it is better at smaller values of $\omega/T$, and for the local spin susceptibility. The better scaling of the local susceptibility is as expected from the localized nature of the spin fluctuations in the Griffiths region. The value of the exponent $\alpha$ decreases significantly with increasing $\lambda$, down to $\alpha\approx 0.5$ at $\lambda=-0.43$. A separate, and in principle, exact determination of $\alpha$ is obtained from the imaginary frequency susceptibility ($\Omega_{m}$) of the quantum Monte results of Ref. Patel et al. (2024b) (Fig. 4E). This also obeys $\Omega_{m}/T$ scaling and yields $\alpha\approx 0.2$ in the Griffiths region, compatible with the our observations at $x=0.22$. We note that $\omega/T$ scaling fails in our theory at very low $T$ Sup , as it does in the Ising model in a transverse field Motrunich et al. (2000), presumably due to the dominance of rare regions: future observations at even lower $T$ are therefore of interest.
The same theory is able to produce a linear resistivity and Planckian behavior in the transport and optical scattering rates Patel et al. (2024a) yielding a consistent picture of the strange metal, as shown in Fig. 4F, with similar results at other values of $\lambda$ in the Supplementary Sup . There is linear-in-$T$ behavior, but with different slopes at lower and higher $T$: we can identify these with the ‘foot’ and ‘fan’ regimes, associated with Griffiths and marginal Fermi liquid behaviors respectively Cooper et al. (2009); Michon et al. (2023); Sachdev (2025). The low $T$ upturn in Fig. 4C in $\kappa$ for $\omega>T$ is possibly also related to the crossover between these ‘foot’ and ‘fan’ regimes.
In summary, we have observed critical spin fluctuations with $\omega/T$ scaling in an overdoped cuprate superconductor. When combined with earlier measurements Aeppli et al. (1997) this shows that there is an extended region of critical spin-fluctuations in La2-xSrxCuO4. We have compared our observations with numerical studies of a Yukawa-Sachdev-Ye-Kitaev model for the onset of spin density wave order in a disordered metal and found good agreement. Both experiment and theory/numerics find an extended ‘Griffiths’ phase of both $\omega/T$ magnetic susceptibility scaling and $T$-linear resistivity. Such scaling is present even though there is no criticality in the usual sense of a diverging correlation length Sup . The correlation length diverges at the lower doping boundary of the Griffiths region, and the large correlation length allows definition of a finite dynamic exponent $z$ which is doping (tuning parameter) dependent. The spin excitations show $z\approx 2$, or larger, and are consistent with contribution with spin relaxation rate $\hbar\Gamma_{\delta}\approx 3k_{B}T$. The observations may be related to singular charge density fluctuations observed in EELS Guo et al. (2024). The strongly $T$-dependent spin fluctuations coexist with transport properties described by Planckian dissipation, showing the spin degrees of freedom in the presence of spatial disorder are at the heart of strange metal behavior. It would be interesting to study the role of such critical spin fluctuations in high temperature $d$-wave superconductivity Abanov et al. (2001).
S.M.H is grateful to Jörg Schmalian for sharing insights about scaling theory. A.A.P. and S.S. thank Peter Lunts for related collaborations. Neutron beamtime was provided by the ISIS neutron and muon source through proposals RB2220248 and RB2410260. The Flatiron Institute is a division of the Simons Foundation. S.S. was supported by the U.S. National Science Foundation grant No. DMR-2245246.